a) Gregg just turned 22 and wants to have $10,283 saved in 8 years by his 30th birthday. If he currently has $6,000 in a bond fund, how much interest does he need to make this happen?

b) You have saved $5,000 for a used car. If you invest it in a CD (Certificate of Deposit), earning 5 percent compounded annually, how long would it take to buy a car that costs $7,755?

c) Linda needs $10,000 in 4 years to use as a down payment on a house. She has $6830 to invest today. What amount of interest does she need to earn to make this happen?
Answer #1

a) Gregg just turned 22 and wants to have $10,283 saved in 8 years by his 30th birthday. If he currently has $6,000 in a bond fund, how much interest does he need to make this happen?

Answer

If interest rate is assumed to be compounded annually on Bond

Amount = Principle(1+r/100)^n

10283 = 6000(1+r/100)^8

(1+r/100)^8 = 10283/6000

(1+r/100)^8 = 1.71

r=7%

Hence by hit and trial method rate of interest at the rate of 7% will give the desired result

b) You have saved $5,000 for a used car. If you invest it in a CD (Certificate of Deposit), earning 5 percent compounded annually, how long would it take to buy a car that costs $7,755?

Answer:

Interest rate is assumed to be compounded annually

Amount = Principle(1+r/100)^n

7755 = 5000(1+5/100)^n

(1+5/100)^n = 7755/5000

(1+5/100)^n = 1.55

n=9

Hence by hit and trial method rate of interest at the rate of 5% will give the desired result in 10 years

c) Linda needs $10,000 in 4 years to use as a down payment on a house. She has $6830 to invest today. What amount of interest does she need to earn to make this happen?

Answer:

Interest rate is assumed to be compounded annually

Amount = Principle(1+r/100)^n

10000 = 6830(1+r/100)^4

(1+r/100)^4 = 10000/6830

(1+r/100)^4 = 1.46

r=10%

Hence by hit and trial method rate of interest at the rate of 10% will give the desired result in 4 years